OpenStudy (anonymous):

PPPLLLZZZ. HELP!!!!!!!!!! Mr. Buff has done 50 push-ups today. Tonight when he is watching t.v. he is going to do 15 push-ups during every commercial break. P stands for the total push-ups Mr.Buff has done today and c stands for the number of commercial breaks. Write an equation for P in terms of c.

jimthompson5910 (jim_thompson5910):

how far did you get with this?

jimthompson5910 (jim_thompson5910):

it might help to think of it like this has already done 50 commercial break #1: 50 + 15*1 commercial break #2: 50 + 15*2 commercial break #3: 50 + 15*3 etc etc try to generalize this for c commercial breaks

OpenStudy (akashdeepdeb):

The total HAS to include the 50 push ups right? And with each commercial he does 15! So 15 * number of commercial breaks?? Now the total is always the sum? You get the answer! :)

OpenStudy (anonymous):

So if c=2 then the equation would be 50+15(2)=P?

jimthompson5910 (jim_thompson5910):

yep that's correct

OpenStudy (akashdeepdeb):

Good job! :)

jimthompson5910 (jim_thompson5910):

so to go from a particular example of 50+15(2)=P to a more general equation, replace 2 with c to get 50+15c=P and you can rearrange things to get P = 15c + 50

OpenStudy (akashdeepdeb):

And then when you generalize that by putting the actual number of pushups (c) You will get your equation! :)

jimthompson5910 (jim_thompson5910):

which hopefully you can see at this point

jimthompson5910 (jim_thompson5910):

I recommend you use other values of c (like c = 3, c = 4, etc) to find the number of pushups done after 3, 4, ... commercial breaks

OpenStudy (anonymous):

I was only in class one day when they were doing graphs. Do you still write out the equation the same or wats different

jimthompson5910 (jim_thompson5910):

not sure what you mean

jimthompson5910 (jim_thompson5910):

but to graph this, you just need 2 points

OpenStudy (anonymous):

We have to use 0,2,5 & 10 for c and. Figure out P

OpenStudy (anonymous):

Rise over run

jimthompson5910 (jim_thompson5910):

to get any point, you plug in a value of c to get a corresponding value of P eg: when c = 0 P = 15c + 50 P = 15*0 + 50 P = 0 + 50 P = 50 so one point is (0, 50)

jimthompson5910 (jim_thompson5910):

what's another point?

OpenStudy (anonymous):

2, 5, and 10

jimthompson5910 (jim_thompson5910):

when c = 2, what is P?

OpenStudy (anonymous):

30

jimthompson5910 (jim_thompson5910):

not quite

jimthompson5910 (jim_thompson5910):

P = 15c + 50 P = 15*2 + 50 ... replace c with 2 (since c = 2) P = ??

OpenStudy (anonymous):

80

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so (2,80) is another point

jimthompson5910 (jim_thompson5910):

what about c = 5?

jimthompson5910 (jim_thompson5910):

technically you just need 2 points to draw this line, so you could stop at plotting those two

OpenStudy (anonymous):

Thank you for helping explain that better

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

P=15c+50 P=15*5+50 p=125

jimthompson5910 (jim_thompson5910):

so (5,125) is another point on this line

OpenStudy (anonymous):

10,150

jimthompson5910 (jim_thompson5910):

that's a bit off

OpenStudy (anonymous):

200

jimthompson5910 (jim_thompson5910):

you forgot to add on 50

jimthompson5910 (jim_thompson5910):

yep, (10,200) is another point

OpenStudy (anonymous):

I was just about to say that lol

jimthompson5910 (jim_thompson5910):

lol good catch